OpenStudy (anonymous):

5=logbase3(x^2+18) convert to exponential equation,simplify and solve to find two solutions

OpenStudy (anonymous):

put both sides as exoponents to the base 5 3^(5) = 3^logbase3(x^2+18) = 3^(5) = x^(2) + 18

OpenStudy (anonymous):

*put both sides as exoponents to the base 3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Now you can just solve for x as the log_3 is elimenated

OpenStudy (anonymous):

\[log_b(x)=y\iff b^y=x\] you should be able to switch back and forth easily. in this case you get \[3^5=x^2+18\]